# Doubly-Dynamic ISAC Precoding for Vehicular Networks: A Constrained Deep Reinforcement Learning (CDRL) Approach

### v2

@@ Abstract
Integrated sensing and communication (ISAC) technology is essential for enabling vehicular networks. However, the communication channel in this scenario exhibits time-varying characteristics, and the potential targets may move rapidly, creating a doubly-dynamic phenomenon. This nature poses a challenge for real-time precoder design. While optimization-based solutions are widely researched, they are complex and heavily rely on perfect channel-related information, which is impractical in double dynamics. To address this challenge, we propose using constrained deep reinforcement learning (CDRL) to facilitate dynamic updates to the ISAC precoder. Additionally, the primal dual-deep deterministic policy gradient (PD-DDPG) and Wolpertinger architecture are tailored to efficiently train the algorithm under complex constraints and varying numbers of users. The proposed scheme not only adapts to the dynamics based on observations but also leverages environmental information to enhance performance and reduce complexity. Its superiority over existing candidates has been validated through experiments.
@@
integrated sensing and communication (ISAC), deep reinforcement learning (DRL), precoder design, double dynamics.
@@ I Introduction
The Internet of Vehicles (IoV) is essential for supporting an efficient and reliable intelligent transportation system, with high-speed communications and high-precision sensing serving as cornerstones [1, 2]. In comparison to separate communication and sensing systems, integrated sensing and communication (ISAC) offers the potential for mutual enhancement, making it a recent research focus in vehicular networks [3, 4, 5, 6]. However, the dynamic nature of the road environment introduces time-varying characteristics in the wireless channel due to user and scatterer movements. Simultaneously, the targets themselves move rapidly and unpredictably, presenting a doubly-dynamic challenge for ISAC.
While previous studies by [7, 8] and [9, 10] have considered ISAC under individual communication or sensing dynamics, implementing ISAC under double dynamics remains unexplored. Compared to single dynamics, the presence of double dynamics introduces multiple challenges to the design of ISAC precoding. Firstly, existing optimization-based methods, e.g., [3, 4], exhibit high computational complexity, resulting in unaffordable latency. Secondly, Doppler shifts in time-varying channels can induce severe inter-carrier interference (ICI) in wideband systems, leading to degraded spectrum efficiency (SE) and sensing accuracy. Moreover, applying optimization-based methods typically requires accurate prior information, including instantaneous channel state information (CSI) and real-time target’s position prior, which can be prohibitively challenging to acquire in double dynamics.
@@ II-A Transmitted Waveform
II-B Time-varying Communication Model
@@
II-C Time-varying Sensing Model
@@ IV Constrained DRL-based Precoding for ISAC without perfect prior
Since perfect prior can hardly be acquired in strong dynamic scenarios, in this section, we introduce a CDRL-based approach for ISAC precoding. The problem will be formulated as a CMDP with appropriate learning strategies for efficient training.
@@ IV-B3 Primal-Dual Network Update
In the learning architecture of the actor and critics, convolutional neural networks (CNN) are employed for processing $\bm{S}_{\text{H}}$ and $\bm{S}_{\text{P}}$ respectively, then the features are concatenated followed by fully-connected networks (FCN). Denote $L_{\text{c}}$ as the maximal number of convolutional layers, $N_{\text{f}}$ as the maximal number of input and output feature maps, and $N_{\text{k}}$ as the maximal side length of the filters in CNN. $N_{\text{L}}$ and $N_{\text{w}}$ as the maximal number of neurons in the hidden layers and the number of hidden layers in FCN. Then the complexity in the forward pass of the actor network in online deployment is $\mathcal{O}(L_{\text{c}}(UG_{\text{t}}N_{\text{d}}+N_{\text{x}}N_{\text{y}})N_{\text{k}}^{2}N_{\text{f}}^{2}+N_{\text{w}}^{2}N_{\text{L}})$, which is linear to $U$ and independent with $N_{\text{t}}$, $M$.

### v3

@@ Abstract
Integrated sensing and communication (ISAC) technology is essential for supporting vehicular networks. However, the communication channel in this scenario exhibits time variations, and the potential targets may move rapidly, resulting in double dynamics. This nature poses a challenge for real-time precoder design. While optimization-based solutions are widely researched, they are complex and heavily rely on perfect channel-related information, which is impractical in double dynamics. To address this challenge, we propose using constrained deep reinforcement learning to facilitate dynamic updates to the ISAC precoder. Additionally, the primal dual-deep deterministic policy gradient and Wolpertinger architecture are tailored to efficiently train the algorithm under complex constraints and varying numbers of users. The proposed scheme not only adapts to the dynamics based on observations but also leverages environmental information to enhance performance and reduce complexity. Its superiority over existing candidates has been validated through experiments.
@@
integrated sensing and communication, deep reinforcement learning, precoder design, double dynamics.
@@ I Introduction
The Internet of Vehicles (IoV) is essential for supporting an efficient and reliable intelligent transportation system, with high-speed communications and high-precision sensing serving as cornerstones [1, 2]. In comparison to separate communication and sensing systems, integrated sensing and communication (ISAC) offers the potential for mutual enhancement, making it a recent research focus in vehicular networks [3, 4, 5, 6, 7]. However, the dynamic nature of the road environment introduces time-varying characteristics in the wireless channel due to user and scatterer movements. Simultaneously, the targets themselves move rapidly and unpredictably, presenting a doubly-dynamic challenge for ISAC.
While previous studies by [8, 9] and [10] have considered ISAC under individual communication or sensing dynamics, implementing ISAC under double dynamics remains unexplored. Compared to single dynamics, the presence of double dynamics introduces multiple challenges to the design of ISAC precoding. Firstly, existing optimization-based methods, e.g., [3, 5], exhibit high computational complexity, resulting in unaffordable latency. Secondly, Doppler shifts in time-varying channels can induce severe inter-carrier interference (ICI) in wideband systems, leading to degraded spectrum efficiency (SE) and sensing accuracy. Moreover, applying optimization-based methods typically requires accurate prior information, including instantaneous channel state information (CSI) and real-time target’s position prior, which can be prohibitively challenging to acquire in double dynamics.
@@ II-A Transmitted Waveform
II-B Communication Model
@@
II-C Sensing Model
@@ IV Constrained DRL-based Precoding for ISAC without perfect prior
Since perfect prior can hardly be acquired in strong dynamic scenarios, henceforth, we introduce a CDRL-based approach for ISAC precoding. The problem will be formulated as a CMDP with appropriate learning strategies for efficient training.
@@ IV-B3 Primal-Dual Network Update
In the learning architecture of the actor and critics, convolutional neural networks (CNN) are employed for processing $\bm{S}_{\text{H}}$ and $\bm{S}_{\text{P}}$ respectively, then the features are concatenated followed by fully-connected networks (FCN). Denote $L_{\text{c}}$ as the maximal number of convolutional layers, $N_{\text{f}}$ as the maximal number of input and output feature maps, and $N_{\text{k}}$ as the maximal side length of the filters in CNN. $N_{\text{L}}$ and $N_{\text{w}}$ as the maximal number of neurons in the hidden layers and the number of hidden layers in FCN. Then the complexity in the forward pass of the actor network in online deployment is $\mathcal{O}(L_{\text{c}}(UG_{\text{t}}N_{\text{d}}+N_{\text{x}}N_{\text{y}})N_{\text{k}}^{2}N_{\text{f}}^{2}+N_{\text{w}}^{2}N_{\text{L}})$.